OpenStudy (anonymous):

The radius r of the nucleus of an atom of mass number A is r = 1.2A1/3 femtometers. Find A if r = 3.6 femtometers.

OpenStudy (anonymous):

solve for A A = 3r/1.2 plug in r = 3.6

OpenStudy (anonymous):

A = 3(3.6)/1.2 A = 10.8/1.2 A = 9

OpenStudy (anonymous):

thats an integer value which is good for a mass number. looks correct